• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name yez

    Log entry time 19:08:22 on February 17, 2011

    Entry number 347379


    keyword=L-HRS: S1 issue updated and some change(Move the cable of S1 TDC from 80~96 to 0~15)

    Today I was trying to fix issue of S1, which has un-correlated ADC and
    TDC singal in replayed data. Here is what I did:
    1, Checked cable connection.   All correct
    2, Checked timing. Both ADC and TDC signals are in the gate (and STOP)
    3, Looked at TDC and ADC signals in the scope. When turn off the HVs,
    both disappear at the same time. Connection is good.
    4, Turned off several PMTs' HV and took cosmic run. Did not see signals
    of ADC and TDC for those channles that has HVs off. So the map in DB is GOOD.
    5, Using a different cable for S1 TDC and move it to different channels
    (from 80~96 to 0~15). Data Bas is updated in 20110217/db_L.s1.dat.Cosmic
    Ray data is taking and will analyze it soo.
    6, Analyzer:
      We guess there may be an error during data replay, since S1, S2m were
    using 1875 TDC and now we changed to 1877. Although we verified the
    cratemap and data base, which are updated, we may have to look carefuly
    on whether the analyzer really treat it as 1877. David will analyze some
    old data we took last week which inlucde S2m trigger and see whether S2m
    has TDC&ADC un-correlated issue. If yes, the data are safe and we need to
    work on analyzer. If not, S1 may has issues on itself and we need to fix
    it asap. 

    A copy of this log entry has been emailed to: camsonne@jlab.org,vasulk@jlab.org, danez@jlab.og